Project K costs $75,000, its expected cash inflows are $8,000 per year for 8 years, and its WACC is 12%. What is the project's MIRR? Round your answer to two decimal places.
Answer #1

MIRR is similar to IRR. However in IRR, it is assumed that intermediary cash flows are reinvested at IRR and in MIRR, it is assumed that intermediary cash flows are reinvsted at WACC.

Thus 75000 has become 98397.55 in eight years

98397.55 = 75000 * ( 1 + MIRR) ^ 8

(1 + MIRR ) ^ 8 = 98397.55 / 75000

= 1.3120

1 + MIRR = 1.3120 ^ (1/8)

= 1.0345

MIRR = 1.0345 - 1

= 0.0345 i.e 3.45%

Pls comment if any further asisstance is required.